OpenStudy (anonymous):

Find the velocity and speed. Find any times which the particle stops: x=cos3t y=sin5t

OpenStudy (anonymous):

sorry that is of no help

OpenStudy (anonymous):

ok sorry that it didn't. Since i didn't help i will give you a metal.

OpenStudy (cwrw238):

x = cos3t dx/dt = -3sin3t dy/dt = 5cos5t dy/dx = -5cos5t/3sin3t velocity at time t = -5cos5t/3sin3t when particle stops -5cos5t/3sin3t = 0

OpenStudy (anonymous):

umm what wld the speed be?

OpenStudy (cwrw238):

not sure of that

OpenStudy (experimentx):

dy/dt is your vy and dx/dt is your vx your speed will be sqrt(vx^2+vy^2)

OpenStudy (anonymous):

ya i know that but i am not sure how to simplify it

OpenStudy (cwrw238):

yes experiment - i forgot that bit

OpenStudy (experimentx):

your speed is going to function of time, try to simplify as much as possible.

OpenStudy (experimentx):

\[\tan\theta=dy/dx\] that will give your the direction, and theta will also be the function of time.

OpenStudy (anonymous):

ok thanks